The circle x^2+y^2=4 cuts the circle x^2+y^2+2x+3y-5=0 in A&B.Then the equation of circle on AB as diameter is?
I tried this by family of circles but didn't got the answer. Please help

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
by family of circle
 
S1 + a S2 = 0
 
x^2 (1 + a) + y^2 (1+a) + 2ax + 3ay – 5a – 4 = 0
 
x^2 + y^2  + 2ax/(1+a) + 3ay/ (1+a) –( 5a + 4)/(1+a) = 0
 
 
now s1 – S2 = 0
 
2x + 3y – 1 = 0
 
now centre of circle passes thriugh this line then you will get a